### v2

# A Sober Look at Progress in Language Model Reasoning: Pitfalls and Paths to Reproducibility

Abstract
Reasoning has emerged as the next major frontier for language models (LMs), with rapid advances from both academic and industrial labs. However, this progress often outpaces methodological rigor, with many evaluations relying on benchmarking practices that lack transparency, robustness, or statistical grounding. In this work, we conduct a comprehensive empirical study and find that current mathematical reasoning benchmarks are highly sensitive to subtle implementation choices—including decoding parameters, random seeds, prompt formatting, and even hardware and software configurations. Performance gains reported in recent studies frequently hinge on unclear comparisons or unreported sources of variance. To address these issues, we propose a standardized evaluation framework with clearly defined best practices and reporting standards. Using this framework, we reassess recent methods and find that most reinforcement learning (RL) approaches yield only modest improvements—far below prior claims—and are prone to overfitting, especially on small-scale benchmarks like AIME’24. In contrast, supervised finetuning (SFT) methods show consistently stronger generalization in the settings we study. To foster reproducibility, we release all code, prompts, and model outputs, for reasoning benchmarks, establishing more rigorous foundations for future work.
1 Introduction
“The first principle is that you must not fool yourself, and you are the easiest person to fool.”
—Richard Feynman
Reasoning has become central to recent advances in large language models (LLMs), playing a key role in nearly all frontier systems (Jaech et al., 2024; Anthropic, 2025; OpenAI, 2025a; xAI, 2025; Meta-AI, 2025; DeepMind, 2025). Recent months have seen a surge of research focused on understanding and improving LLM reasoning, accompanied by several open-source tools and training strategies (see Li et al. (2025c) for a survey). This momentum has sparked optimism that building capable, competitive reasoning models may soon be within reach.
However, as evaluation practices shape the direction and perceived progress of the field (Liao et al., 2021), concerns around methodological rigor are growing. Non-reproducible or inconclusive evaluations can distort scientific understanding, misguide adoption, and skew future research priorities (Henderson et al., 2018; Marie et al., 2021; Musgrave et al., 2020; Prabhu et al., 2020; Andrychowicz et al., 2020; Colas et al., 2018). In the fast-moving area of LLM reasoning, where rapid publication cycles and benchmarking races are common, methodological shortcuts can quietly undermine progress. While concerns about reproducibility in LLM evaluations are well-documented (Reuel et al., 2024; Biderman et al., 2024), their persistence - especially in reasoning, calls for renewed scrutiny and higher standards.
Motivated by a growing number of inconsistent empirical claims across the reasoning landscape, we conduct a rigorous investigation into the current state of reasoning benchmarks, focusing specifically on mathematical reasoning - one of the most widely used testbeds for evaluating algorithmic advances in this space (HuggingFaceH4, 2024; AI-MO, 2024a).
Our main finding is that many recent empirical conclusions may be overly optimistic and fail to generalize under careful re-evaluation. We identify a surprising degree of sensitivity in LLM-based reasoning pipelines to seemingly minor design choices, ranging from decoding parameters, prompt formatting, and random seeds, to the hardware and software stacks used during evaluation (see Table 1). Particularly concerning is the instability introduced by small benchmark sizes: for example, AIME’24 and AMC’23 each contain only 30–40 examples, making performance metrics highly volatile, where even one question can shift Pass@1 by over 3 percentage points. This leads to substantial variance across seeds, often resulting in double-digit performance swings that challenge the reliability of published results. In Section 3, we systematically analyze the root causes of this instability, including sampling variance, decoding configurations, evaluation frameworks, and hardware heterogeneity. We show that these factors can significantly distort conclusions if not carefully controlled.
In Section 4, we propose a set of best practices aimed at improving reproducibility and rigor in reasoning benchmarks. We also re-evaluate recent techniques using a standardized and reproducible evaluation stack. Our findings are sobering: reinforcement learning (RL) applied to distillation-based models such as DeepSeek-R1 yields little to no statistically significant gains. Some methods, such as OpenRS, show promising results in original reports, but fail to hold up under repeated evaluation. RL training on base models like Qwen2.5 Math does show stronger performance, but still often underperforms instruction-tuned counterparts.111We note that OpenReasoner-Zero is a consistent exception, achieving competitive performance. Furthermore, RL-trained models exhibit significant performance drops on newer benchmarks such as AIME’25, echoing patterns of test set overfitting or “hill-climbing” observed in prior work (Golchin & Surdeanu, 2023; Roberts et al., 2023; Dominguez-Olmedo et al., 2024). In contrast, supervised fine-tuning (SFT) continues to deliver stable, generalizable improvements across benchmarks, underscoring its maturity as a training paradigm. These observations point to a critical need for more reliable and standardized evaluation protocols.
Taken together, in this work, we aim to provide not only a clearer assessment of where current methods stand, but also the tools and practices needed to make reasoning evaluation more transparent, robust, and reproducible. To this end, we open-source all code, prompts, and outputs to facilitate fair and accountable progress in this increasingly important area.
| Model | Algorithm | Base | Framework | Temp | Top_p | Seq. Len | K |
| DeepSeek-R1-Distill-1.5B | SFT | Qwen2.5-Math-1.5B | – | 0.6 | 0.95 | 32,768 | 64 |
| DeepSeek-R1-Distill-7B | SFT | Qwen2.5-Math-7B | – | 0.6 | 0.95 | 32,768 | 64 |
| DeepSeek-R1-Distill-14B | SFT | Qwen2.5-14B | – | 0.6 | 0.95 | 32,768 | 64 |
| DeepSeek-R1-Distill-32B | SFT | Qwen2.5-32B | – | 0.6 | 0.95 | 32,768 | 64 |
| OpenThinker-32B | SFT | Qwen2.5-32B-Instruct | evalchemy | 0.7 | 0.8 | 32,768 | 5 |
| Bespoke-Stratos-32B | SFT | Qwen2.5-32B-Instruct | evalchemy | 0.7 | 0.8 | 32,768 | 5 |
| Bespoke-Stratos-7B | SFT | Qwen2.5-7B-Instruct | evalchemy | 0.7 | 0.8 | 32,768 | 5 |
| s1.1-7B | SFT | Qwen2.5-7B-Instruct | lm-eval-harness | 0 | – | 32,768 | 64 |
| s1.1-32B | SFT | Qwen2.5-32B-Instruct | lm-eval-harness | 0 | – | 32,768 | 64 |
| LIMO | SFT | Qwen2.5-32B-Instruct | math-eval-harness | 0 | 1 | 32,768 | 1 |
| MiniMath-R1-1.5B | SFT | DeepSeek-R1-Distill-1.5B | oumi-ai | – | – | – | – |
| Light-R1-7B | SFT | DeepSeek-R1-Distill-7B | verl | 0.6 | 0.95 | 32,768 | – |
| DeepScaleR-1.5B-Preview | RL | DeepSeek-R1-Distill-1.5B | verl | 0.6 | 0.95 | 32,768 | 16 |
| L1-Exact | RL | DeepSeek-R1-Distill-1.5B | verl | 0.6 | 0.95 | 8,000 | – |
| L1-Max | RL | DeepSeek-R1-Distill-1.5B | verl | 0.6 | 0.95 | 8,000 | – |
| Open-RS1 | RL | DeepSeek-R1-Distill-1.5B | lighteval | 0.6 | 0.95 | 32,768 | 32 |
| Open-RS2 | RL | DeepSeek-R1-Distill-1.5B | lighteval | 0.6 | 0.95 | 32,768 | 32 |
| Open-RS3 | RL | DeepSeek-R1-Distill-1.5B | lighteval | 0.6 | 0.95 | 32,768 | 32 |
| II-Thought-1.5B-Preview | RL | DeepSeek-R1-Distill-1.5B | evalscope | 0.6 | 0.95 | 32,768 | 64 |
| Oat-Zero-1.5B | RL | Qwen2.5-Math-1.5B | custom | 0 | 1 | 3,000 | 1 |
| Oat-Zero-7B | RL | Qwen2.5-Math-7B | custom | 0 | 1 | 3,000 | 1 |
| STILL-3-1.5B-preview | RL | DeepSeek-R1-Distill-1.5B | custom | 0.6 | 0.95 | 32,768 | 5 |
| FastCurl-1.5B-Preview | RL | DeepSeek-R1-Distill-1.5B | verl | 0.6 | 0.95 | 32,768 | 16 |
| LIMR | RL | Qwen2.5-Math-7B | custom | 0.4 | 0.95 | 3,072 | 4 |
| SimpleRL-Zoo-7B | RL | Qwen2.5-7B | verl | 1 | 0.95 | 16,000 | 32 |
| SimpleRL-Zoo-Math-7B | RL | Qwen2.5-Math-7B | verl | 1 | 0.95 | 16,000 | 32 |
| OpenReasoner-Zero-7B | RL | Qwen2.7-7B | – | 1 | 1 | – | 16 |
| Eurus2 Prime | RL | Qwen2.5-Math-7B | custom | 0 | 1 | 4,096 | – |
2 Related Works
Language Model Reasoning (for Math). The recent releases of OpenAI-o1 (Jaech et al., 2024) (in September 2024), OpenAI-o3 (OpenAI, 2025a) (in December 2024) and DeepSeek-R1 (DeepSeek-AI, 2025) (in January 2025), have spurred the language modelling community to work on improving the reasoning capabilites of language models. Several popular methods for improving those capabilites have emerged with supervised fine-tuning (SFT) and reinforcement learning (RL) being the two primary methods of interest (Uesato et al., 2022; Lightman et al., 2023; Lyu et al., 2025; Open Thoughts, 2025). Recent works have built upon the DeepSeek-R1 recipe by proposing newer RL algorithms, including LCPO (Aggarwal & Welleck, 2025), REINFORCE++ (Hu, 2025), DAPO (Yu et al., 2025), DPO-VP (Tu et al., 2025), VinePPO (Kazemnejad et al., 2024), CPPO (Lin et al., 2025a), VAPO (Yue et al., 2025b) and GRO (Cai, 2025). To gain a stronger understanding of how to induce mathematical capabilities, other works have conducted significant empirical studies exploring the design space of RL methods (Zeng et al., 2025b; Liu et al., 2025c; Team et al., 2025; Shao et al., 2024), including data scaling trends (Shen et al., 2025), curriculums (Wen et al., 2025; Roux et al., 2025) and reward design (Gao et al., 2024a; Cui et al., 2025; Ma et al., 2023). Based on the success of these methods, there have also been recent efforts into scaling up reinforcement learning approaches to induce reasoning in domains beyond math, including code (Liu & Zhang, 2025; Xie et al., 2025; Jha et al., 2024; Yu et al., 2024), medicine (Zhang et al., 2025; Sim & Chen, 2024) and other sciences (Su et al., 2025; Yuan et al., 2025; Zeng et al., 2025a). Further, some works also explored scaling up RL-based approaches to modalities beyond just language, including vision (Ma et al., 2025; Meng et al., 2025; Huang et al., 2025; Peng et al., 2025; Chen et al., 2025; Deng et al., 2025; Liu et al., 2025d; Feng et al., 2025; Lin et al., 2025b). In our work, we objectively re-evaluate the claims made by several of these recent works under a standardized lens, and find that many of the reported gains do not hold up strongly when pitted on a level-playing field against well-tuned baselines.
Sobering Studies on ML Progress. Machine learning is a field of rapid progress. Due to the lightning speed of papers coming out across the various sub-fields of machine learning, practitioners and researchers often fail to rigorously evaluate algorithmic progress (Hutchinson et al., 2022; Dehghani et al., 2021; Machado et al., 2018; Ghosh et al., 2024; Balduzzi et al., 2018; Liao et al., 2021; Cawley & Talbot, 2010; Lipton & Steinhardt, 2019; Prabhu et al., 2024b; Card et al., 2020; Dror et al., 2018). This has led to several papers showing that simple well-tuned baselines outperform months of progress on a specific sub-field in machine learning, including in continual learning (Prabhu et al., 2024a; 2020), active learning (Cawley, 2011) and test-time adaptation (Press et al., 2023). With the rapid influx of reasoning-based LMs, such statistically rigorous comparisons of models are ever more important—yet, despite the heavy use of RL-algorithms for driving progress in reasoning, there is very little mention of how different methods standardize their evaluations across different factors of variability. RL-algorithms themselves are known to be quite fickle to extremely minor variations including random seeds (Agarwal et al., 2021; Gorsane et al., 2022; Chan et al., 2019; Jordan et al., 2020; Patterson et al., 2024). Some works have even gone as far as suggesting that reliable benchmarking of RL-based methods is computationally infeasible (Jordan et al., 2024). Additionally, other works have demonstrated critical reliability issues in the generalization of frontier models to minor perturbations in the question inputs (Mirzadeh et al., 2024; Nezhurina et al., 2024; Srivastava et al., 2024), the type of tasks tested (Yan et al., 2025b; Petrov et al., 2025; Dominguez-Olmedo et al., 2024; Roberts et al., 2025), metrics used (Liu et al., 2024) and in data-scarce scenarios (Udandarao et al., 2024; Kandpal et al., 2023; Parashar et al., 2024). Given such a volatile landscape, in this work, we aim to level the playing field across recent LM-methods that have been released and provide an objective look on the progress that the reasoning community has made. Our findings, which we discuss in the rest of the paper, are sobering at best.
3 Exploring the Design Space of Reasoning: What Matters Most?
Recent reasoning focused language models are evaluated under highly heterogeneous conditions—including differences in evaluation frameworks, hardware, number of random seeds, temperature, and nucleus sampling parameters (top_p) (see Table 1). While prior work has examined the effect of sampling parameters in multiple-choice (Renze, 2024) and coding tasks (Arora et al., 2024), the influence of these choices remains underexplored for open-ended reasoning models—particularly those trained with reinforcement learning. In this section, we systematically assess how these evaluation design choices affect reported performance, and highlight the sources of variance that most impact the reliability of results.
3.1 Experimental Setup
We adopt a consistent experimental setup throughout this section, unless otherwise stated. Our analysis includes nine widely used models grouped into two commonly benchmarked size classes: 1.5B and 7B parameters. For the 1.5B class, we evaluate: DeepSeek-R1-Distill-1.5B (DeepSeek-AI, 2025), DeepScaleR-1.5B (Luo et al., 2025), II-1.5B-Preview (Intelligent Internet, 2025) , OpenRS1-1.5B, OpenRS2-1.5B, and OpenRS3-1.5B (Dang & Ngo, 2025). Note that DeepScaleR-1.5B, II-1.5B-Preview, and the OpenRS models are all initialized from DeepSeek-R1-Distill-1.5B and subsequently finetuned via reinforcement learning (e.g., GRPO (Shao et al., 2024)) to enhance mathematical reasoning capabilities. For the 7B class, we evaluate: DeepSeek-R1-Distill-7B, S1.1-7B (Muennighoff et al., 2025), and OpenThinker-7B (Open Thoughts, 2025). Both S1.1-7B and OpenThinker-7B are finetuned Qwen2.5-7B-Instruct models (Qwen et al., 2025), trained using supervised learning on reasoning traces derived from DeepSeek-R1. All models are benchmarked on three widely used datasets: AIME’24 (AI-MO, 2024a), AMC’23 (AI-MO, 2024b), and MATH500 (Hendrycks et al., 2021), using the Pass@1 metric. Each result is averaged over multiple seeds and obtained on a standardized software stack (through a Docker image), and hardware with the following configuration: one 40 GB A100 GPU, an AMD 7302 32-core CPU, and 1TB RAM. All experiments were run using lighteval (Fourrier et al., 2023) with the vllm backend (Kwon et al., 2023).
Sampling Parameters: To systematically compare the impact of sampling parameters on accuracy, all experiments in this section were performed with a standardized configuration: temperature=0.8, top_p=0.9, and both max_model_len and max_new_tokens set to 32,768 tokens. This context length matches the limits of models such as OpenThinker-7B and S1.1-7B, although certain models (e.g., DeepSeek) support longer sequences of up to 131,072 tokens. We chose this standardized evaluation length to ensure comparability, with a detailed analysis of the influence of completion length presented in Figure 8. Unless otherwise specified, results in this section are averaged over 10 random seeds for AIME’24 and AMC’23, and 3 seeds for MATH500, following the recommendations from Section 3.2.1.
3.2 Seed Variance in Evaluation
We begin by analyzing the variance induced purely by the random seed used during evaluation—an aspect often neglected in benchmarking practices. While recent work calls for statistical rigor (e.g., using error bars and multiple runs) (Bowyer et al., 2025; Biderman et al., 2024; Madaan et al., 2024), evaluations frequently rely on single-seed runs, obscuring potential variability. We assess the seed-induced variance across 20 independent evaluation runs for each of the nine models. Results are shown in Figure 2.
Key Insight. Pass@1 values show surprisingly high standard deviation—ranging from 5 to 15 percentage points across seeds. This issue is particularly severe for AIME’24 and AMC’23, which have only 30 and 40 test samples respectively. A change in just one question shifts Pass@1 by 2.5–3.3 percentage points.
3.2.1 Can Bootstrapping Improve Mean Estimates?
To mitigate high variance, recent work has adopted bootstrapping: averaging multiple evaluation runs to stabilize results. For example, DeepSeek reports Pass@1 over 64 runs, while DeepScaleR uses 16. We study the effectiveness of this approach by bootstrapping estimates for AIME’24 using 1 to 60 evaluation runs. Figure 3 shows that while variance is extreme for and still large for , it reduces sharply for . Further analysis of variance across additional datasets is presented in Figures 14 and 16.
3.2.2 Variance from Sampling Parameters: Temperature and top-p
We additionally investigate the impact of the temperature and top_p hyperparameter as prior works often employ different temperature and top_p settings when comparing the same model. To isolate the impact of varying temperature and top_p, we averaged pass@1 across seeds and compute variation of this estimate across temperature and top_p in a boxplot. Figure 4 and 5 show the performance variation. We see that temperature-induced and top_p-induced fluctuations not only affect performance estimates but also introduce substantial variability in performance itself, which can lead to unfair comparisons when evaluating the same model across different temperatures.
Reducing the temperature or increasing the nucleus sampling parameter (top_p) improves the accuracy of performance estimates without incurring additional computational cost. Figure 6 illustrate the impact of temperature and Figure 7 show that of top_p across multiple models and datasets. Notably, a more reproducible estimate is associated with significant drops in measured performance, highlighting a consistent tradeoff between reproducibility and high performance. We recommend optimizing the temperature for performance, and comparing the best parameter per model.
3.3 Effect of Prompt Format and Context Length
Maximum Output Tokens. Figure 8 shows that reducing max_new_tokens harms performance—especially on long-form problems. This sensitivity varies by model and dataset. Although reducing this setting lowers cost, it may induce premature stopping, leading to incorrect answers.
Prompt Format. Prompt formatting has a measurable impact on accuracy. As shown in Figure 9, models perform best when using math-specific prompts and their native chat templates. Omitting templates leads to performance drops, particularly for instruction-tuned models. We compare accuracy under three different prompt settings (see Table 6): (1) a math-specific prompt formatted using the model’s chat template, (2) only the model’s chat template with no additional prompt, and (3) no template at all, i.e., the question without any special tokens or instructions. Interestingly, while base models like Qwen2.5-Math may benefit from prompt-free setups (Liu et al., 2025c), instruction-tuned models rely heavily on format alignment. Thus, maintaining consistent and format-aware prompting is essential for maximizing instruction-tuned model performance.
3.4 Variance from Hardware & Software Factors: A Critical Consideration
Performance can also vary due to non-obvious factors like hardware and evaluation framework—yet this is rarely acknowledged. Models are often tested on heterogeneous systems and evaluated using different toolchains. For example, S-1.1 (Muennighoff et al., 2025) uses lm-evaluation-harness (Gao et al., 2024b), the OpenRS model suite uses lighteval (Fourrier et al., 2023), and II-1.5B-Preview uses evalscope (Alibaba ModelScope Community, 2025) for evaluation.
Hardware Variation. We evaluated the same model across five different compute clusters, each with varying GPU types and memory configurations. As shown in Figure 10, performance varied by up to 8% for OpenRS-1.5B and 6% for DeepSeek-R1-Distill-7B on AIME’24, with similar trends observed on AMC’23. While it is known that inference engines such as vLLM can be sensitive to hardware differences (vLLM Contributors, 2024)—and that low-level optimizations in PyTorch or CUDA (PyTorch Contributors, 2024) may introduce non-determinism—our results demonstrate that these effects can measurably impact benchmark accuracy, even when averaging over multiple seeds.
Mitigation strategies. We conducted extensive experiments to isolate and quantify hardware-induced variability, even when using identical GPU types and software stacks. Our investigation had 3 stages, each attempting to eliminate additional sources of variation:
Stage 1: Standardized Docker across different A100 clusters. We first evaluated models using our standardized Docker container on A100 GPUs across two platforms: our internal cluster and Runpod (a cloud GPU service). Despite identical hardware specifications and containerized environments, we observed notable performance differences. For instance, Qwen2.5-Math-1.5B achieved 7.3%3.8 on AIME’24 on our cluster versus 11.3%3.6 on Runpod; a 4 percentage point gap that exceeds the standard deviation.
Stage 2: Enforcing CUDA determinism. Suspecting non-deterministic GPU operations, we enforced strict determinism by setting torch.use_deterministic_algorithms(True), fixing CUDA seeds, and disabling cuDNN benchmarking. Comparing A100 and H100 clusters, variance decreased but persisted: OpenThinker2-7B scored 53.0%4.6 on A100 versus 57.1%5.2 on H100 for AIME’24. Even with deterministic algorithms, hardware architectural differences and vLLM’s backend variations (vLLM Contributors, 2024) continued to introduce discrepancies.
Stage 3: Identical hardware, updated stack. Finally, we updated all components (vLLM, LightEval, math-verify) and ran multiple evaluations on identical hardware. Remarkably, even consecutive runs on the same A100 cluster produced variations: Bespoke-Stratos-7B ranged from 19.3% to 23.0% on AIME’24 across runs. These persistent differences, despite controlling for hardware, software versions, and random seeds highlight deep-seated sources of non-determinism in modern inference stacks, potentially arising from dynamic kernel selection, memory allocation patterns, or floating-point accumulation order (Atil et al., 2025).
As shown in Figure 10, these effects are not merely theoretical but measurably impact benchmark accuracy even when averaging over multiple seeds. This underscores that reproducibility challenges extend beyond simple seed variance to fundamental architectural and systems-level factors (Atil et al., 2025; Yao et al., 2025; He & Lab, 2025).
Evaluation across different Python frameworks. Evaluation results can vary based on the framework used, due to differences in prompt templates, inference engines (e.g., vLLM (Kwon et al., 2023)), and response extraction strategies (e.g., MathVerify). For example: lighteval is used by OpenRS (Dang & Ngo, 2025), evalchemy (Raoof et al., 2025) is used by models like OpenThinker and Bespoke-Stratos, other frameworks include lm-evaluation-harness (Gao et al., 2024b) and evalscope (Alibaba ModelScope Community, 2025). Some prior works have also studied the performance differences induced by using different evaluation frameworks for evaluating base language models (Li et al., 2024).
To assess this impact, we compare lighteval and evalchemy, fixing all other variables: model, dataset, hardware, decoding parameters, and random seeds (3 per model). For a fair comparison, we evaluated two models, DeepSeek-R1-Distill-1.5B and S1.1-7B, using default temperature and top_p parameter values on a single GPU. We present results averaged over three seeds for higher robustness. As shown in Table 2, framework-induced differences are generally small (1–2pp) but can still affect model rankings in tightly clustered scenarios.
| Model | lighteval | evalchemy |
| R1-Distill-1.5B | 26.6 | 26.6 |
| S1.1-7B | 22.2 | 17.7 |
Overall, our findings underscore that significant performance variations can arise solely from differences in hardware and software configurations, emphasizing the need to standardize for reliable evaluations.
4 Way Forward: Standardization in Evaluations
In this section, we standardize evaluation frameworks, propose best practices, and comprehensively evaluate existing methods.
4.1 Recommendations: Which practices to adopt?
We propose a set of best practices informed by our experiments and guided with current research insights:
Hardware and Software Stack Standardization: To promote reproducibility and facilitate future work, we release all code within a Docker container, along with step-by-step instructions for running experiments on Runpod’s publicly accessible, on-demand GPU instances. This setup allows any researcher to replicate and extend our results under identical conditions.
Variance Estimates: For small benchmarks (e.g., AIME’24), run evaluations with at least 30 random seeds. Report the mean and standard deviation to quantify uncertainty and assess the statistical significance of performance differences.
Model-Specific Hyperparameter Optimization: Tune hyperparameters (such as temperature and top_p) separately for each model, then fix them across tasks to ensure consistency and fair comparisons.
Context Length and Prompt Template Selection: Ensure the context length is sufficiently large, especially for models with long reasoning chains to avoid premature truncation and under-reported accuracy. For instruction-tuned models, always use the appropriate chat template to match the expected input format.
Robust Answer Extraction: We strongly recommend using a resilient answer extraction pipeline that handles parsing issues and evaluates expression equivalence, rather than relying on exact string matching. This reduces the likelihood of spurious gains from formatting artifacts.
Transparent Evaluation Protocols: We recommend to release code, prompts, and model outputs, and clearly document the evaluation stack. Report uncertainties (e.g., via standard deviations) and include both quantitative and qualitative analyses to enable thorough and reproducible comparisons.
4.2 Standardization Procedure
We adopt a largely consistent experimental setup with prior work, with the key difference being our use of publicly accessible cloud instances from Runpod222https://www.runpod.io/pricing. Each instance is equipped with a single A100 PCIe GPU, 8 vCPUs, and 128 GB of RAM. We evaluate all models listed in Table 3 across six benchmarks: AIME’24 (AI-MO, 2024a), AIME’25 (Lin, 2025), AMC’23 (Knovel Engineering, 2025), MATH500 (HuggingFaceH4, 2024), Minerva (Lewkowycz et al., 2022), and OlympiadBench (He et al., 2024). All experiments are conducted using the LightEval framework (Fourrier et al., 2023) (0.8.1) with a vLLM backend, repeated across ten random seeds for AIME’24, AIME’25, AMC’23 and three random seeds for the rest. While our analysis in section 3.2.1 identifies 30 seeds as optimal for variance stabilization, we use 10 seeds for AIME/AMC benchmarks as a practical compromise between rigor and computational cost. This still represents a significant improvement over the single-seed evaluations common in prior work, and we report standard deviations to quantify remaining uncertainty. Depending on the base model architecture, we set the maximum number of new tokens (e.g., 4096 for QwenMath-based models), apply optimal hyperparameters, and use the standardized chat template except for base models. LightEval’s LaTeX-based answer extraction and evaluation pipeline ensures reliable and consistent result parsing and correctness matching, similar to math-verify.
Statistical Significance Testing. To rigorously assess whether RL-trained models genuinely outperform their baselines, we employ paired sample-wise statistical tests rather than simply comparing mean accuracies. For each model-baseline pair, we: (1) compute the average accuracy per problem instance across all random seeds, (2) align these sample-wise accuracies between the RL/SFT model and the baseline from which it was trained, and (3) conduct one-sided paired tests to evaluate whether the RL model achieves systematically higher performance. We employ both parametric (paired t-test) and non-parametric (Wilcoxon signed-rank) tests to ensure robustness to distributional assumptions. The null hypothesis is that there is no improvement from RL/SFT training, with the alternative hypothesis that the RL/SFT model performs better than its baseline. We report results as statistically significant at two thresholds: (marked with ∗) and (marked with ∗∗). This conservative approach ensures that reported improvements represent genuine performance gains rather than statistical artifacts from multiple testing or random variation.
4.3 A Sober Look: Results
| Model | AIME’24 | AIME’25 | AMC’23 | MATH500 | Minerva | Olympiad |
| Based on: Deepseek R1 Distill Qwen 1.5B (RL) | ||||||
| R1-Distill (DeepSeek-AI, 2025) | 28.74.8 | 22.35.2 | 71.53.9 | 84.90.3 | 30.51.0 | 52.40.4 |
| L1-Exact (Aggarwal & Welleck, 2025) | 24.43.3 | 22.34.2 | 70.53.7 | 86.60.8 | 31.51.7 | 52.51.3 |
| L1-Max (Aggarwal & Welleck, 2025) | 27.74.2 | 21.05.0 | 73.26.0 | 84.70.1 | 33.30.9 | 52.30.6 |
| Open-RS1 (Dang & Ngo, 2025) | 28.96.0 | 21.34.2 | 75.03.3 | 85.10.8 | 30.40.2 | 53.21.9 |
| Open-RS2 (Dang & Ngo, 2025) | 31.37.7 | 22.75.6 | 73.05.7 | 84.10.2 | 29.21.1 | 53.70.6 |
| Open-RS3 (Dang & Ngo, 2025) | 29.74.6 | 24.76.5 | 69.25.5 | 84.21.1 | 28.62.3 | 51.80.8 |
| STILL-3 (Min et al., 2024) | 34.75.5 | 24.06.4 | 72.55.4 | 86.61.9 | 30.00.6 | 53.91.5 |
| ZR1-1.5B (Zyphra, 2025) | 30.34.6 | 24.03.8 | 78.84.0 | 86.60.9 | 32.10.9 | 56.40.4∗∗ |
| II-Thought (Intelligent Internet, 2025) | 32.05.9 | 24.04.1 | 79.55.1∗ | 86.60.6 | 31.70.6 | 54.90.4∗ |
| DeepScaleR (Luo et al., 2025) | 37.06.6 | 30.34.3∗∗ | 76.24.6 | 87.81.0∗∗ | 31.01.5 | 55.51.1∗∗ |
| FastCuRL (Song et al., 2025) | 36.34.6∗ | 26.53.7 | 76.94.8 | 87.51.0∗∗ | 30.81.6 | 56.70.8∗∗ |
| DisCO-logL (Li et al., 2025a) | 39.04.5 | 31.74.8∗ | 78.83.4∗ | 87.41.1∗ | 32.71.3 | 58.60.3∗∗ |
| Nemotron-RR (Liu et al., 2025a) | 48.36.3∗∗ | 33.05.1∗ | 87.04.8∗∗ | 91.00.5∗∗ | 37.51.6∗∗ | 62.80.9∗∗ |
| Based on: Deepseek R1 Distill Qwen 7B (RL) | ||||||
| R1-Distill (DeepSeek-AI, 2025) | 52.36.3 | 39.05.9 | 91.52.7 | 94.10.3 | 40.10.4 | 67.30.1 |
| Sky-T1 (mini) (Team, 2025a) | 51.34.5 | 39.05.7 | 90.03.3 | 93.31.0 | 41.81.2 | 67.70.9 |
| DisCO-logL (Li et al., 2025a) | 51.33.9 | 41.34.5 | 92.22.2 | 94.01.3 | 44.71.1∗∗ | 67.60.3 |
| Skywork-OR1 (He et al., 2025) | 60.74.7∗ | 49.76.2∗ | 94.03.8 | 95.70.9∗ | 43.01.0∗ | 73.60.9 ∗∗ |
| Based on: Deepseek R1 Distill Qwen 7B (SFT) | ||||||
| R1-Distill (DeepSeek-AI, 2025) | 52.36.3 | 39.05.9 | 91.52.7 | 94.10.3 | 40.10.4 | 67.30.1 |
| Light-R1 (Wen et al., 2025) | 53.04.8 | 41.03.5 | 90.03.1 | 93.50.5 | 41.31.3 | 68.01.2 |
| Based on: Qwen2.5 Math 1.5B (RL) | ||||||
| Math (Base) (Yang et al., 2024) | 11.33.6 | 5.72.7 | 44.04.9 | 51.75.5 | 11.32.2 | 26.00.6 |
| Oat-Zero (Liu et al., 2025b) | 16.03.2 | 6.73.4 | 52.52.9 | 73.51.7∗∗ | 26.30.8∗∗ | 37.21.3∗∗ |
| Math (Instruct) (Yang et al., 2024) | 12.01.7 | 11.75.7∗ | 54.85.3 | 74.70.5∗∗ | 26.71.8∗∗ | 37.90.2∗∗ |
| LUFFY-Zero (Yan et al., 2025a) | 15.94.9 | 11.72.8 | 54.55.4 | 77.61.2∗∗ | 27.20.4∗∗ | 42.30.5∗∗ |
| Based on: Qwen2.5 Math 7B (RL) | ||||||
| Math (Base) (Yang et al., 2024) | 20.73.8 | 8.73.9 | 56.25.7 | 64.30.5 | 17.31.9 | 29.00.5 |
| SimpleRL-Zoo (Zeng et al., 2025b) | 22.75.2 | 10.73.4 | 62.23.6 | 76.91.8∗∗ | 30.12.8∗∗ | 39.30.6∗∗ |
| LIMR (Li et al., 2025b) | 30.73.2 | 7.83.3 | 62.23.4 | 76.50.4∗∗ | 34.91.3∗∗ | 39.30.9∗∗ |
| Oat-Zero (Liu et al., 2025b) | 28.03.1 | 8.82.5 | 66.23.6 | 79.40.3∗∗ | 34.41.4∗∗ | 43.81.1∗∗ |
| Sky-T1 (Team, 2025a) | 19.32.6 | 21.03.9∗ | 67.23.6∗∗ | 85.20.5∗∗ | 34.71.1∗∗ | 52.10.8∗∗ |
| Math (Instruct) (Yang et al., 2024) | 15.73.9 | 10.73.8 | 67.03.9 | 82.90.1∗∗ | 35.00.6∗∗ | 41.30.9∗∗ |
| LUFFY-Zero (Yan et al., 2025a) | 26.76.5 | 23.35.7∗ | 73.53.9∗∗ | 87.90.1∗∗ | 34.12.2∗∗ | 54.90.3∗∗ |
| Based on: Qwen2.5 1.5B (RL) | ||||||
| Qwen (Base) (Qwen et al., 2025) | 0.00.0 | 0.00.0 | 2.52.5 | 3.31.5 | 1.80.4 | 1.50.5 |
| SimpleRL-Zoo (Zeng et al., 2025b) | 0.31.1 | 0.31.1 | 13.24.7∗ | 12.06.5∗∗ | 4.02.4∗ | 4.22.0∗∗ |
| Qwen (Instruct) (Qwen et al., 2025) | 1.31.7 | 0.71.4 | 26.24.8∗∗ | 58.11.4∗∗ | 19.41.3∗∗ | 20.40.5∗∗ |
| Based on: Qwen2.5 7B (RL) | ||||||
| Qwen (Base) (Qwen et al., 2025) | 8.03.2 | 3.73.7 | 35.84.9 | 59.70.3 | 21.41.9 | 27.01.0 |
| SimpleRL-Zoo (Zeng et al., 2025b) | 14.02.1 | 4.32.7 | 58.01.6∗∗ | 77.90.8∗∗ | 33.00.2∗∗ | 39.00.1∗∗ |
| Open Reasoner Zero (Hu et al., 2025) | 19.72.9 | 15.72.7 | 59.54.5∗∗ | 83.91.1∗∗ | 31.61.3∗∗ | 47.61.7∗∗ |
| Qwen (Instruct) (Qwen et al., 2025) | 12.33.2 | 7.33.4 | 52.84.8∗∗ | 77.30.8∗∗ | 34.91.0∗∗ | 38.90.9∗∗ |
| Based on: Qwen2.5 7B Instruct (SFT) | ||||||
| Qwen (Instruct) (Qwen et al., 2025) | 12.33.2 | 7.33.4 | 52.84.8 | 77.30.8 | 34.91.0 | 38.90.9 |
| Eurus2 Prime (Cui et al., 2025) | 18.71.7 | 14.31.6 | 64.84.0 | 80.10.1∗∗ | 37.51.0∗∗ | 43.90.3∗∗ |
| s1.1 (Muennighoff et al., 2025) | 19.03.2 | 21.05.5∗ | 59.53.7 | 80.80.6∗ | 37.51.1 | 48.21.4∗∗ |
| Bespoke Stratos (Bespoke Labs, 2024) | 20.34.3 | 18.04.8 | 60.24.9 | 84.70.5∗∗ | 39.11.3∗ | 51.91.1∗∗ |
| OpenThinker (Open Thoughts, 2025) | 30.56.2∗∗ | 26.04.4∗∗ | 71.43.9∗∗ | 88.31.4∗∗ | 37.93.8∗∗ | 55.61.4∗∗ |
| OpenR1 (Hugging Face, 2025) | 48.38.9∗∗ | 35.54.2∗∗ | 86.04.5∗∗ | 93.50.7∗∗ | 41.21.3∗∗ | 67.41.3∗∗ |
| OpenThinker2 (Open Thoughts, 2025) | 53.04.6∗∗ | 41.05.0∗∗ | 87.03.5∗∗ | 94.40.7∗∗ | 42.01.5∗∗ | 70.61.0∗∗ |
| OpenThinker3 (Open Thoughts, 2025) | 66.75.2∗∗ | 57.06.4∗∗ | 91.82.6∗∗ | 95.90.6∗∗ | 42.80.8∗∗ | 74.80.8∗∗ |
| Model | AIME’24 | AIME’25 | AMC’23 | MATH500 | Minerva | Olympiad |
| Based on: Qwen2.5 32B (RL) | ||||||
| Qwen (Base) (Qwen et al., 2025) | 8.05.5 | 2.32.2 | 36.56.8 | 62.83.5 | 28.92.2 | 29.52.0 |
| DAPO (Yu et al., 2025) | 43.04.0∗∗ | 32.36.1∗∗ | 88.53.6∗∗ | 91.00.3∗∗ | 40.00.6∗∗ | 60.91.1∗∗ |
| Qwen (Instruct) (Qwen et al., 2025) | 15.74.7 | 13.04.6 | 60.85.7 | 81.60.5 | 40.31.3 | 46.71.3 |
| Based on: Qwen QwQ 32B (RL) | ||||||
| QwQ-32B (Team, 2025b) | 76.33.3 | 69.04.5 | 96.22.4 | 97.50.6 | 49.00.2 | 78.11.0 |
| INTELLECT-2 (Prime Intellect, 2025) | 75.22.4 | 66.36.6 | 96.02.7 | 97.00.3 | 49.80.2 | 78.00.8 |
| Based on: DeepSeek R1 Distill Qwen 32B (SFT) | ||||||
| Distill R1 (DeepSeek-AI, 2025) | 67.01.9 | 55.35.7 | 96.82.1 | 95.10.7 | 45.11.7 | 73.80.5 |
| TinyR1-Preview (Sun et al., 2025) | 73.34.4 | 66.06.6∗ | 97.82.2 | 97.20.4∗∗ | 48.50.5∗ | 76.31.2∗ |
| Light-R1 DS (Wen et al., 2025) | 76.74.7∗ | 68.75.7∗∗ | 97.01.6 | 97.30.1∗∗ | 48.21.1∗ | 76.90.8∗∗ |
| Based on: Qwen2.5 32B Instruct (SFT) | ||||||
| Qwen (Instruct) (Qwen et al., 2025) | 15.74.7 | 13.04.6 | 60.85.7 | 81.60.5 | 40.31.3 | 46.71.3 |
| Sky-T1-Preview (Team, 2025a) | 37.08.1∗∗ | 22.73.1 | 75.83.5∗∗ | 88.60.5∗∗ | 39.81.2 | 56.30.8∗∗ |
| Bespoke-Stratos (Bespoke Labs, 2024) | 37.75.9∗∗ | 26.37.3∗ | 80.73.9∗∗ | 90.30.3∗∗ | 43.31.5 | 61.51.7∗∗ |
| LIMO (Ye et al., 2025) | 57.84.7∗∗ | 46.35.5∗∗ | 93.02.8∗∗ | 94.61.2∗∗ | 45.22.3∗ | 70.50.9∗∗ |
| s1.1-32B (Muennighoff et al., 2025) | 61.36.9∗∗ | 49.06.9∗∗ | 94.23.9∗∗ | 94.80.5∗∗ | 46.20.8∗∗ | 72.40.7∗∗ |
| OpenThinker (Open Thoughts, 2025) | 68.56.7∗∗ | 50.36.2∗∗ | 95.02.9∗∗ | 95.50.3∗∗ | 46.81.2∗∗ | 71.80.5∗∗ |
| OpenThinker2 (Open Thoughts, 2025) | 71.36.5∗∗ | 61.75.3∗∗ | 95.82.9∗∗ | 96.10.4∗∗ | 46.60.6∗∗ | 75.50.6∗∗ |
| Standalone Models (No tools) | ||||||
| GPT-OSS-20B (medium) (OpenAI, 2025b) | 75.75.0 | 72.75.4 | 97.03.3 | 96.50.8 | 45.11.1 | 77.00.7 |
We present experimental results in Table 3, and analyze different aspects of the results.
RL-training on R1-Distill. We evaluated several reinforcement learning (RL) approaches (e.g., GRPO) using the DeepSeek R1-Distill-1.5B model. We first observe that none of the L1 models (Aggarwal & Welleck, 2025) outperformed the original DeepSeek R1-Distill baseline — an expected outcome given that L1 training prioritized smaller output length over accuracy. OpenRS (Dang & Ngo, 2025) reported strong gains (10–15%) on AIME, AMC, and OlympiadBench. However, our replication showed no statistically significant improvements over the R1-Distill baseline. Same case held for Still-3 and Light-R1 models, which showed no significant improvement over the R1-Distill baseline. II-Thought yields modest improvements across benchmarks, especially over AIME’24 but the observed gains did not carry over significantly to AIME’25 indicating overfitting to existing benchmarks. Only DeepscaleR and FastCuRL demonstrated statistically significant improvements across many benchmarks. Notably, recent models like Nemotron-RR deliver the first recipes with robust improvements over the DeepSeek R1-Distill baseline.
RL Training on Qwen2.5 Math and Base Models. We next analyze RL training applied to the Qwen2.5 Base and Qwen2.5 Math Base models, a trend trying to replicate gains by Deepseek-R1 Zero. Unlike the R1-Distill results, RL training with Oat-Zero, LIMR, and SimpleRL-Zoo consistently produced statistically significant gains over the base model, especially across Math500, Minerva and OlympiadBench benchmarks. This indicates that RL-based approaches can indeed offer substantial improvements given a base model instead of a distilled R1 model. However, these gains are often roughly comparable and sometimes slightly higher than achieved via instruction tuning in the original Qwen papers, suggesting that instruction tuning with additional math data alone may be sufficient to far achieve the current gains from RL methods in this setting. We also observed that the improvements on AIME’24 were also significant, but did not carry over to AIME’25 indicating a troubling overfitting trend.
Effectiveness of Supervised Finetuning. We assessed supervised finetuning methods like s1.1, Eurus2 Prime, Bespoke Stratos, OpenR1 and OpenThinker models, which further refine instruction-tuned models using reasoning traces. Supervised methods consistently outperformed the instruct-tuned baseline across all benchmarks (even Minerva) and generalized comparatively well to AIME’25. The performance improvements from OpenThinker series and OpenR1 were especially notable, showcasing the promise of data curation for supervised finetuning (Guha et al., 2025). These results underscore the maturity and effectiveness of SFT when training recipes are scaled to large datasets.
Scaling to Larger Parameters. We extend our analysis to 32B scale models in Table 4. Among R1-Distill variants, Light-R1 DS achieves significant performance improvements, demonstrating that supervised finetuning can further improve even strong distilled models. For Qwen2.5-32B base models, SFT approaches like OpenThinker series achieve strong gains (71.3% and 68.5% on AIME’24 respectively) over the instruct model, further emphasising the importance of high-quality data curation. Notably, DAPO - an RL approach, shows improvements over the base model, and showcases strong performance gains even relative to the Qwen Instruct model. The QwQ-32B models demonstrate state-of-the-art performance (76.3% on AIME’24), and INTELLECT-2 shows no meaningful improvement over QwQ-32B despite additional training. These results reinforce our earlier findings: (1) SFT on reasoning traces remains the most reliable approach for performance gains, and (2) scaling to larger models amplifies these benefits, with 32B models achieving substantially higher absolute performance than their smaller counterparts.
Overfitting and Generalization. We now examine the overfitting by comparing performance on AIME’24 versus the more challenging AIME’25. RL-trained models showed a pronounced performance drop between the two, indicating overfitting to the training distribution. In contrast, supervised fine-tuning (SFT) models maintained consistent improvements, suggesting better generalization. Openthinker2 showed significant degradation compared to Openthinker across benchmarks not provided in their blogpost, indicating overfitting via data-curation. This highlights a gap in current evaluation protocols, and a need to assess out-of-distribution generalization for reasoning models.
Having established these standardized benchmarking results, we now examine whether several recently reported phenomena in reasoning models replicate under our rigorous evaluation framework.
5 Do Discovered Phenomena Replicate? A Detailed Analysis.
We further investigate three recently noted phenomena to see if they replicate in our experiments: (1) how response length correlates with performance, (2) the decline in response diversity following reasoning-focused training and (3) the effect of spurious prompts on performance
5.1 Are Incorrect Responses Longer?
Recent research (Wang et al., 2025) suggests that incorrect answers often have disproportionately long reasoning chains. We first verify whether this finding holds in our setting, and then we explore possible explanations behind the observed variations.
Do longer responses indicate a higher likelihood of an incorrect answer? We compare the distribution of response lengths for correct and incorrect answers across 6 datasets (AIME24, AIME25, AMC23, MATH500, Minerva and OlympiadBench) averaged across random seeds for each model. Figure 11 shows histograms of the average number of responses per seed, binned by response length. A clear trend emerges: shorter responses are significantly more likely to be correct, while longer responses become progressively more error-prone. This pattern is consistent across all seeds and is especially pronounced for responses exceeding 10,000 tokens. We now address two questions:
Q1. Does this pattern hold for both RL- and SFT-trained models?
Yes. We find the trend is consistent across both RL- and SFT-trained models (see Appendix figures 22 and 23 for detailed per-model breakdown ). We consistently observe that the effect is more pronounced in RL-trained models (displayed on the left) than in SFT-trained models (displayed on the right). As detailed in the Appendix, both the Qwen 2.5 Math base exhibit a slight shift in length, though this shift is notably more evident in R1-distill and subsequent RL-trained models.
Q2. Is this primarily because of truncated or incomplete responses?
No, truncation is not the primary cause. While responses approaching the 32,000-token limit are almost always incorrect, the correlation between length and errors persists even for complete responses well below this limit (e.g., 10,000-15,000 tokens). This indicates that excessive length itself, rather than truncation, is associated with reasoning failures.
5.2 Is There Diversity Collapse in Reasoning Training?
Dang et al. (2025) and Yue et al. (2025a) have reported a counterintuitive phenomenon in reasoning models: improvements in Pass@1 achieved through supervised fine-tuning or RL can reduce Pass@k performance due to diminished output diversity—a phenomenon termed diversity collapse. Theoretical analyses attribute this collapse to the model concentrating too much probability mass on a single reasoning path, while current decoding strategies fail to recover the lost diversity.
To examine these claims, we compare the Pass@k performance (for ) of two RL-trained models (DeepScaleR-1.5B and FastCuRL-1.5B) against their base model DeepSeek-R1-Distill-Qwen-1.5B across all datasets. Figures 12 and 24 show the delta in Pass@k relative to DeepSeek-R1-Distill-Qwen-1.5B.
Findings. We do observe a minor diversity collapse. Gains in Pass@1 generally come with regression in Pass@k, though the magnitude of the decay varies.
5.3 Spurious Prompts and Qwen2.5-Math-7B
[…]
6 Conclusion
Our study shows that much of the perceived progress in LLM-based reasoning, particularly in mathematical benchmarks, rests on unstable and often non-reproducible foundations. We find that minor differences in sampling parameters, prompt formatting, hardware, and software configurations can lead to major shifts in reported performance—casting doubt on many recent empirical claims. Reinforcement learning methods, while promising in theory, offer at best modest gains in practice and are prone to overfitting, especially on small benchmarks like AIME’24. In contrast, supervised finetuning continues to deliver consistent, generalizable improvements across a wide range of benchmarks and model sizes. To address these challenges, we advocate for standardized, transparent evaluation protocols. Our open-sourced framework, complete with Dockerized environments, seed-averaged metrics, and robust answer matching, provides reproducible foundations for future research. We hope this work shifts the focus from leaderboard chasing to methodological rigor—ensuring that future claims of progress in reasoning are both meaningful and measurable.
Author Contributions
Andreas, Vishaal and Ameya conceived the project. Andreas and Hardik co-led the experiments, with Vishaal and Ameya advising the experimental design. The manuscript was written by Andreas, Hardik, Vishaal and Ameya. Matthias and Samuel provided helpful feedback and advice throughout the project.
Acknowledgments
The authors would like to thank (in alphabetical order): Matteo Farina, Shyamgopal Karthik, Nikhil Parthasarathy, Shiven Sinha, Joschka Strüber, Thaddäus Wiedemer for helpful feedback on the draft. AH acknowledges funding by the Federal Ministry of Education and Research (BMBF), FKZ: 01IS24079A. HB has received funding from the Digital Europe Programme under grant agreement No 101195233 (OpenEuroLLM). AH, HB and VU thank the International Max Planck Research School for Intelligent Systems (IMPRS-IS) for support. VU also thanks the European Laboratory for Learning and Intelligent Systems (ELLIS) PhD program for support. VU was supported by a Google PhD Fellowship in Machine Intelligence. AP and MB acknowledge financial support by the Federal Ministry of Education and Research (BMBF), FKZ: 011524085B and Open Philanthropy Foundation funded by the Good Ventures Foundation. This work was supported by the Digital Europe Programme under grant agreement No 101195233 (OpenEuroLLM).
Appendix
We now provide thorough details about the benchmark, baselines, and prompts.
Appendix A Bootstrapping Results on Additional Datasets
To complement our analysis in Section 3, we present bootstrapped variance results on two additional datasets: AMC’23 and MATH500. As shown in Figures 14 and 15, high variance in Pass@1 persists even when averaging over multiple seeds (), mirroring the trends observed on AIME’24. These results reinforce our conclusion that small benchmark sizes yield unstable estimates and that robust performance reporting requires multiple seed runs. Even for larger Benchmarks, like MATH500 (Figure 16), Minerva (Figure 17) and Olympiad Bench (Figure 18) the estimates remain volative, yet there magnitude is much smaller. An overview of the exact variance values can be found in Table 5.
| Model | |||||||||
| AIME’24 | |||||||||
| DeepSeek-R1-Distill-1.5B | 40.044 | 19.841 | 13.269 | 7.762 | 3.754 | 2.409 | 1.686 | 1.054 | 0.365 |
| DeepScaleR-1.5B | 27.452 | 13.477 | 9.117 | 5.342 | 2.495 | 1.587 | 1.149 | 0.705 | 0.246 |
| FastCuRL-1.5B | 28.807 | 14.647 | 9.488 | 5.658 | 2.711 | 1.709 | 1.247 | 0.757 | 0.256 |
| AIME’25 | |||||||||
| DeepSeek-R1-Distill-1.5B | 24.300 | 11.954 | 7.759 | 4.568 | 2.199 | 1.419 | 1.017 | 0.628 | 0.209 |
| DeepScaleR-1.5B | 17.311 | 8.561 | 5.597 | 3.387 | 1.605 | 0.988 | 0.726 | 0.445 | 0.156 |
| FastCuRL-1.5B | 16.201 | 8.232 | 5.412 | 3.221 | 1.504 | 1.000 | 0.710 | 0.424 | 0.145 |
| AMC’23 | |||||||||
| DeepSeek-R1-Distill-1.5B | 22.544 | 11.241 | 7.301 | 4.288 | 2.093 | 1.315 | 0.942 | 0.585 | 0.202 |
| DeepScaleR-1.5B | 22.377 | 10.851 | 7.167 | 4.228 | 2.079 | 1.295 | 0.936 | 0.568 | 0.192 |
| FastCuRL-1.5B | 23.671 | 11.841 | 7.649 | 4.605 | 2.149 | 1.404 | 0.980 | 0.594 | 0.211 |
| MATH500 | |||||||||
| DeepSeek-R1-Distill-1.5B | 1.175 | 0.561 | 0.367 | 0.226 | 0.110 | 0.069 | 0.049 | 0.030 | 0.010 |
| DeepScaleR-1.5B | 1.050 | 0.513 | 0.351 | 0.198 | 0.099 | 0.062 | 0.044 | 0.027 | 0.009 |
| FastCuRL-1.5B | 0.939 | 0.447 | 0.305 | 0.179 | 0.085 | 0.055 | 0.040 | 0.024 | 0.008 |
| Minerva | |||||||||
| DeepSeek-R1-Distill-1.5B | 3.485 | 1.741 | 1.143 | 0.665 | 0.326 | 0.206 | 0.149 | 0.091 | 0.031 |
| DeepScaleR-1.5B | 2.385 | 1.179 | 0.769 | 0.452 | 0.220 | 0.143 | 0.102 | 0.061 | 0.022 |
| FastCuRL-1.5B | 2.703 | 1.347 | 0.876 | 0.527 | 0.244 | 0.161 | 0.113 | 0.069 | 0.024 |
| Olympiad Bench | |||||||||
| DeepSeek-R1-Distill-1.5B | 1.269 | 0.622 | 0.427 | 0.241 | 0.118 | 0.076 | 0.053 | 0.032 | 0.011 |
| DeepScaleR-1.5B | 1.018 | 0.505 | 0.334 | 0.195 | 0.093 | 0.059 | 0.043 | 0.026 | 0.009 |
| FastCuRL-1.5B | 0.908 | 0.463 | 0.305 | 0.179 | 0.086 | 0.054 | 0.039 | 0.024 | 0.008 |
Appendix B Prompt Variants and Template Settings
We provide the exact templates used for our three prompt settings in Table 6: Math, Default, and No Template. These formats are based on the DeepSeek tokenizer but adapted for each model’s specific chat template. Our results (in Section 3.3) indicate that instruction-tuned models are highly sensitive to prompt formatting, with performance degrading significantly when prompts deviate from their training-time structure.
| Prompt | Example |
[…]
| Prompt | Example |
[…]
Appendix C Hardware Differences
In Figure 19, we show that similar discrepancies are observed on MATH500 as observed in AIME and AMC – showing the variance from hardware is observed even on larger test sets.
Appendix D Effect of Output Length Limits
We further explore how varying max_new_tokens impacts model accuracy. Figures below compare OpenRS-series models (with 131,072-token context windows) and OpenThinker/S1.1 models (with 32,768-token limits).
Figure 20 shows that OpenRS models are highly sensitive to this parameter—shortening outputs results in clear accuracy drops. Similarly, Figure 21 reveals the same pattern for OpenThinker-7B and S1.1-7B, despite their smaller context lengths. In both cases, premature truncation leads to incomplete reasoning chains and incorrect answers, confirming the importance of setting appropriate generation limits.
Appendix E Response Length vs. Accuracy — Per-Model Breakdown
To supplement the aggregated results shown in Figure 11, we include detailed histograms for each individual model in the appendix. These plots show the distribution of correct and incorrect responses across response lengths, averaged over random seeds. Due to the number of models analyzed, we split the results into two figures for clarity.
Figures 22 and 23 reveal that the overall trend observed in the main paper holds consistently across nearly all models: incorrect responses tend to be longer than correct ones.
These results reinforce the idea that excessively long outputs often indicate failure modes such as hallucinated reasoning, verbose overthinking, or degenerate loops. Importantly, this correlation persists well below the maximum sequence length, ruling out truncation as the sole cause. Across all models, longer responses are a consistent marker of incorrect outputs, making response length a useful signal for detecting low-confidence or erroneous reasoning chains.
Appendix F Diversity Collapse
In addition to the diversity collapse discussed in Section 5.2 and Figure 12 the results for additional datasets are shown in Figure 24. For DeepSeek-R1-Distill-1.5B (SFT-trained) we could not replicate a diversity collapse as shown in Figure 25.
Appendix G Optimal Decoding Parameters
Empirically, the decoding parameters in Table 8 consistently produced optimal performance for the models we evaluated.
| Model Name | Temperature | Top-p |
| agentica-org/DeepScaleR-1.5B-Preview | 1.0 | 0.7 |
| bespokelabs/Bespoke-Stratos-7B | 1.0 | 0.9 |
| deepseek-ai/DeepSeek-R1-Distill-Qwen-1.5B | 0.9 | 0.7 |
| deepseek-ai/DeepSeek-R1-Distill-Qwen-7B | 0.9 | 0.7 |
| hkust-nlp/Qwen-2.5-7B-SimpleRL-Zoo | 0.5 | 0.5 |
| Intelligent-Internet/II-Thought-1.5B-Preview | 1.0 | 0.5 |
| knoveleng/Open-RS1 | 1.0 | 0.6 |
| knoveleng/Open-RS2 | 0.9 | 0.8 |
| knoveleng/Open-RS3 | 0.5 | 0.9 |
| l3lab/L1-Qwen-1.5B-Exact | 0.5 | 0.7 |
| l3lab/L1-Qwen-1.5B-Max | 0.7 | 0.9 |
| Nickyang/FastCuRL-1.5B-Preview | 1.0 | 0.7 |
| Open-Reasoner-Zero/Open-Reasoner-Zero-7B | 0.7 | 0.5 |
| open-thoughts/OpenThinker-7B | 0.8 | 0.95 |
| qihoo360/Light-R1-7B-DS | 0.7 | 1.0 |
| RUC-AIBOX/STILL-3-1.5B-preview | 1.0 | 0.6 |
| simplescaling/s1.1-7B | 1.0 | 0.9 |
| GAIR/LIMR | 0.6 | 0.6 |
| open-r1/OpenR1-Qwen-7B | 0.8 | 0.9 |
| Qwen/Qwen2.5-1.5B-Instruct | 0.2 | 1.0 |
| Qwen/Qwen2.5-7B-Instruct | 0.4 | 0.95 |
| Qwen/Qwen2.5-Math-1.5B | 0.7 | 0.5 |
| Qwen/Qwen2.5-Math-7B | 0.5 | 0.5 |
| sail/Qwen2.5-Math-1.5B-Oat-Zero | 0.6 | 0.6 |
| sail/Qwen2.5-Math-7B-Oat-Zero | 0.6 | 0.6 |